OpenStudy (ksaimouli):

y=tan(2x-x^3) u=2x-x^3 find the derivative

OpenStudy (ksaimouli):



OpenStudy (ksaimouli):

AM I RIGHT

OpenStudy (precal):

use www.wolframalpha.com to check yourself

hartnn (hartnn):

d/dx (2x-x^3) =?

OpenStudy (ksaimouli):

2-3x^2

OpenStudy (cwrw238):

yes

hartnn (hartnn):

so whats your answer now ?

OpenStudy (ksaimouli):

but we need to use the chain rule right so 2-3x^2=0-6x again chain rule so 6

OpenStudy (ksaimouli):

i mean -6

hartnn (hartnn):

u do d/dx(2x-x^3) only once, right ? its just a function of x, not a function of function of x so, no chain rule there

OpenStudy (ksaimouli):

what do u mean by function of function of x

hartnn (hartnn):

like sin(x^2) is function of function of x sin x is just function., x^2 is just function chain rule applies to function of function of x

OpenStudy (ksaimouli):

ohk now i get it thx for ur patience and help

hartnn (hartnn):

ok, so the answer here u get is ?

OpenStudy (ksaimouli):

sec^2(2x-x^3))(2-3x^2)

hartnn (hartnn):

correct :)